Could you please help me out with the following question?? Please...
Add
z 6
----- + -----
z^2-3z-10 z+2
Thank you for all your help!

Answer by edjones(8007)

z/(z^2-3z-10) + 6/(z+2)
=z/((z-5)(z+2)) + 6/(z+2)
=(z+6(z-5))/...
=(z+6z-30)/...
=(7z-30)/(z^2-3z-10)
.
Ed